OpenStudy (anonymous):

what is the error in linear approximation when f(x)=sqrt(3+x) a=22 and delta x= .5

OpenStudy (blockcolder):

First of all, what's the linear approximation of your function when a=22?

OpenStudy (anonymous):

5+1/10(x-22)??

OpenStudy (blockcolder):

The error is given by delta(x)*(dy/dx at x=a).

OpenStudy (anonymous):

so it would be (.5)*5?

OpenStudy (anonymous):

no thats not right how do i find derivative at x=a?

OpenStudy (blockcolder):

First, find f'(x), then substitute x=a there. :D

OpenStudy (anonymous):

i got 1/10 for f^1 of (x)

OpenStudy (blockcolder):

Then just multiply that by 0.5 and that's the error. :D

OpenStudy (amistre64):

f(a+.5) - the linear approx of f(a+.5)

OpenStudy (amistre64):

f(x)=sqrt(3+x) a=22 and h = .5 f(a) = sqrt(25) = 5 f'(a) = 1/2sqrt(3+a) = 1/10 y = 1/10(22+.5-22) + 5 y = 5.05 f(25.5) - 5.05 = error